OpenStudy (anonymous):

Please help ^-^ Convert the mixed number to an improper fraction. A. B. C. D.

OpenStudy (compassionate):

Nothing shows.

jabez177 (jabez177):

Download this and take a screenshot and paste the link here. :) http://prntscr.com/9mxt58

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (alexandervonhumboldt2):

\[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\]

OpenStudy (anonymous):

http://prntscr.com/9mxv2c

OpenStudy (alexandervonhumboldt2):

use the rule i gave: we have \[3\frac{ 5 }{ 6 }=\frac{ 3*6+5 }{ 6 }=?\]

OpenStudy (anonymous):

Okay.

jabez177 (jabez177):

Welcome to OpenStudy, Mate! :D